If A and B are independent events, P(A)=0.11 , and P(B)=0.74 , what is P(B|A) ?
Answer #1

For independent events

P(A and B) =P(A)*P(B)

using the formula P(B|A) = P(A and B)/P(A) = [P(A)*P(B)]/P(A) = P(B)..........P(A) cancelled out.

so, P(B|A) = P(B)

we have P(B) = 0.74

therefore, P(B|A) = 0.74